# Mitigating Hallucinations in Large Vision-Language Models via Entity-Centric Multimodal Preference Optimization

### v1

@@ Abstract
Large Visual Language Models (LVLMs) have demonstrated impressive capabilities across multiple tasks. However, their trustworthiness is often challenged by hallucinations, which can be attributed to the modality misalignment and the inherent hallucinations of their underlying Large Language Models (LLMs) backbone. Existing preference alignment methods focus on aligning model responses with human preferences while neglecting image-text modality alignment, resulting in over-reliance on LLMs and hallucinations. In this paper, we propose Entity-centric Multimodal Preference Optimization (EMPO), which achieves enhanced modality alignment than existing human preference alignment methods. Besides, to overcome the scarcity of high-quality multimodal preference data, we utilize open-source instruction datasets to automatically construct high-quality preference data across three aspects: image, instruction, and response. Experiments on two human preference datasets and five multimodal hallucination benchmarks demonstrate the effectiveness of EMPO, e.g., reducing hallucination rates by 85.9% on Object-HalBench and 49.8% on MM-HalBench. The code and dataset will be released at https://github.com/RobitsG/EMPO.
@@ 1 Introduction
Large Vision Language Models (LVLMs) have recently demonstrated impressive capabilities in multimodal question answering Chen et al. (2023); Liu et al. (2023b); Bai et al. (2023); Lu et al. (2024), which typically consists of a visual encoder to extract image features, and a Large Language Model (LLM) to answer the image-related textual instructions based on the provided visual context. The LVLMs are usually learned in two steps Li et al. (2023a); Du et al. (2022); Lin et al. (2024): (1) pretraining on large-scale image-text pairs to learn multimodal knowledge, and (2) fine-tuning on multimodal instruction-following datasets to steer their responsiveness to user instructions Liu et al. (2023b); Wang et al. (2024c); Bai et al. (2023).
Figure 1: Two types of hallucinations. a) Modality misalignment: LVLM recognizes the presence of entities but confuses their semantics. b) LLM inherent hallucination: The LVLM’s response is entirely dependent on textual context, disregarding the image content.
@@ 2 Related Work
Recent research on LVLMs Zhu et al. (2023); Liu et al. (2023b); Bai et al. (2023); Lu et al. (2024); Zhang et al. (2024); Li et al. (2023a) constructs LVLMs by aligning LLMs with visual models, demonstrating superior performance across various visual–language tasks compared to earlier studies Jia et al. (2021); Radford et al. (2021); Ju et al. (2022); Alayrac et al. (2022). These LVLMs typically adopt a two-stage training strategy: (1) Pretraining on large-scale image–text pairs to learn fundamental multimodal knowledge Li et al. (2023a); Du et al. (2022); Lin et al. (2024); Bai et al. (2023), and (2) Instruction fine-tuning using instruction datasets to improve instruction-following abilities Chen et al. (2024b); Wang et al. (2024c); Bai et al. (2023); Wang et al. (2024a); Li et al. (2023a, 2024). For instance, LLaVA Li et al. (2024) introduces synthetic instructions to fine-tune an instruction-following LVLM, while MiniGPT-v2 Chen et al. (2023) employs unique task identifiers during fine-tuning to reduce instruction ambiguity.
@@ 3.1 Preliminaries
Direct Preference Optimization (DPO) Rafailov et al. (2024) is a primary method for human preference alignment that implicitly models the reward function in Reinforcement Learning from Human Feedback Yu et al. (2024b). Given an instruction $q$, chosen response $y_{w}$, and rejected response $y_{l}$, DPO formulates the reward function as
@@ 3.3 Entity-centric Multimodal Preference Optimization
[…]
@@
where the LVLMs are optimized to fully align hallucination-related key facts with human preference, reducing its hallucinations.
@@ 4 Experiments
  $\text{CHAIR}_{s}$downarrow   $\text{CHAIR}_{i}$downarrow   Hall. $\downarrow$  Score $\uparrow$   Acc. $\uparrow$  F1 $\uparrow$
@@
DPO (RLAIF-V DataSet) LLaVA-1.5 7B 19.13 9.32 36.6 2.70 76.8 81.5
@@ A.1 Quality Control
To rigorously assess the effectiveness of our entity-level image editing strategy, we conducted a detailed analysis using CLIP-based semantic similarity metrics. For each edit, we separately evaluated the similarity between the image region and both the original (source) entity label as well as the intended (target) entity label, before and after editing. This dual-perspective evaluation enables objective measurement of whether the edit not only increases the semantic presence of the desired entity, but also reduces residual evidence of the original entity. We employed CLIP Jiang et al. (2023) to compute the semantic similarity between localized image regions and corresponding entity labels. Figure 6(a) overlays the distributions of CLIP similarity scores for both the original and target entities before and after editing. The results show a pronounced and desirable shift: after editing, the similarity between the image region and the target entity rises sharply, forming a high-density mode near 0.8, while the similarity to the original entity drops toward zero. Before editing, the pattern is reversed, with image regions matching the original entity but not the intended target. This demonstrates our method’s ability to both install semantic cues for the target entity and erase clues of the original entity. Entity-wise statistics (Figure 6(b)) show that this effect is robust across all five entity types. Regardless of the category, the target entity’s median similarity rises after editing, while the original entity’s similarity markedly falls, confirming entity token manipulation at the vision-language level. Categories such as “person” and “car” benefit most, reflecting their strong visual signatures in both CLIP training and entity editing efficacy. To quantify the impact, Figure 6(c) presents the empirical distribution of similarity changes. Target entity similarity typically improves by about 0.6 on average post-editing, while the original entity’s similarity most often decreases, with minimal overlap in the two distributions, indicating reliable entity substitution while minimizing unintended blending or hallucination. Finally, Figure 6(d) summarizes success rates across a range of similarity thresholds. For thresholds above 0.5, over 90% of image regions now surpass the success criterion for the target entity, while the corresponding rate for the original entity drops near zero. This robust separation directly validates that our method transforms image evidence is both semantically meaningful and discriminative.
In summary, our CLIP-based analysis verifies that the proposed editing pipeline achieves two critical goals: (1) enhancing alignment between edited image regions and their newly assigned (target) entities, and (2) effectively suppressing residual evidence of the original entities. This fine-grained control supports our wider claim that both image and language data in our preference construction exhibit clear, contrastive, and controllable entity-centric differences while maintaining structural consistency.
@@ A.2 Cost analysis
Having established robust quality safeguards, we now consider quantifying the computational requirements and monetary costs of each processing step. Extracting image entities costs about \$0.002 per entry; generating all 17,332 image entries on 8 A100 GPUs takes roughly 4–6 hours. Instruction-level entity extraction and rewriting add \$0.002 and \$0.003 per entry, respectively. In total, the labeling cost is \$0.007 per entry plus 8.3 seconds of compute time on a single A100. By contrast, manual annotation would cost approximately \$0.30 per entry, over 40× higher.
@@ Appendix B Hallucination Cause Analysis Experiment
These findings highlight the challenges LVLMs face in achieving robust multimodal understanding and highlight the need for improved mechanisms to mitigate hallucinations. Addressing these issues is essential for enhancing the reliability and applicability of LVLMs in real-world tasks that require both textual and visual reasoning.
Appendix C Prompt Appendix
The section describes the GPT4o-mini prompt for identifying entities, as well as the prompts for rewriting chosen instructions and rejected instructions.
@@ Appendix C Prompt Appendix
Provide only the rewritten question without any additional explanation.’’’
@@ 4. Output Format:
Appendix D Example Appendix
@@
Figure 7: Who in the picture are wearing trousers?

### v2

@@ Abstract
Large Visual Language Models (LVLMs) have demonstrated impressive capabilities across multiple tasks. However, their trustworthiness is often challenged by hallucinations, which can be attributed to the modality misalignment and the inherent hallucinations of their underlying Large Language Models (LLMs) backbone. Existing preference alignment methods focus on aligning model responses with human preferences while neglecting image-text modality alignment, resulting in over-reliance on LLMs and hallucinations. In this paper, we propose Entity-centric Multimodal Preference Optimization (EMPO), which achieves enhanced modality alignment compared to existing human preference alignment methods. Besides, to overcome the scarcity of high-quality multimodal preference data, we utilize open-source instruction datasets to automatically construct high-quality preference data across three aspects: image, instruction, and response. Experiments on two human preference datasets and five multimodal hallucination benchmarks demonstrate the effectiveness of EMPO, e.g., reducing hallucination rates by 85.9% on Object-HalBench and 49.8% on MM-HalBench. The code and dataset will be released at https://github.com/RobitsG/EMPO.
@@ 1 Introduction
Large Vision Language Models (LVLMs) have recently demonstrated impressive capabilities in multimodal question answering Chen et al. (2023); Liu et al. (2023b); Bai et al. (2023); Lu et al. (2024); Lyu et al. (2025), which typically consists of a visual encoder to extract image features, and a Large Language Model (LLM) to answer the image-related textual instructions based on the provided visual context. The LVLMs are usually learned in two steps Li et al. (2023a); Du et al. (2022); Lin et al. (2024): (1) pretraining on large-scale image-text pairs to learn multimodal knowledge, and (2) fine-tuning on multimodal instruction-following datasets to steer their responsiveness to user instructions Liu et al. (2023b); Wang et al. (2024c); Bai et al. (2023).
Figure 1: Two types of hallucinations. a) Modality misalignment: LVLM recognizes the presence of entities but confuses their semantics (e.g., misinterpreting a sign as "No Parking" instead of "No Motor Vehicles Allowed"). b) LLM inherent hallucination: The LVLM’s response is entirely dependent on textual context, disregarding the image content (e.g., inferring a "car" must be present whenever a "road" is present).
@@ 2 Related Work
Recent research on LVLMs Zhu et al. (2023); Liu et al. (2023b); Bai et al. (2023); Lu et al. (2024); Zhang et al. (2024); Xu et al. (2025) constructs LVLMs by aligning LLMs with visual models, demonstrating superior performance across various visual–language tasks compared to earlier studies Jia et al. (2021); Radford et al. (2021); Ju et al. (2022); Alayrac et al. (2022). These LVLMs typically adopt a two-stage training strategy: (1) Pretraining on large-scale image–text pairs to learn fundamental multimodal knowledge Li et al. (2023a); Du et al. (2022); Lin et al. (2024); Bai et al. (2023), and (2) Instruction fine-tuning using instruction datasets to improve instruction-following abilities Chen et al. (2024b); Wang et al. (2024c); Bai et al. (2023); Wang et al. (2024a); Li et al. (2023a, 2024). For instance, LLaVA Li et al. (2024) introduces synthetic instructions to fine-tune an instruction-following LVLM, while MiniGPT-v2 Chen et al. (2023) employs unique task identifiers during fine-tuning to reduce instruction ambiguity.
@@ 3.1 Preliminaries
Direct Preference Optimization (DPO) Rafailov et al. (2024) is a primary method for human preference alignment that implicitly models the reward function in Reinforcement Learning from Human Feedback Yu et al. (2024b). By training on preference pairs that distinguish between accurate and hallucinated responses, DPO reduces the model’s tendency to produce hallucinations. Given an instruction $q$, chosen response $y_{w}$, and rejected response $y_{l}$, DPO formulates the reward function as:
@@ 3.3 Entity-centric Multimodal Preference Optimization
$\$- $\beta\log\frac{\pi_{\theta}\left(y\mid v,q_{l}\right)}{\pi_{\mathrm{ref}}\left(y\mid v,q_{l}\right)}),$
@@
where the LVLMs are optimized to fully align hallucination-related key facts with human preference, reducing their hallucinations.
@@ 4 Experiments
  $\text{CHAIR}_{s}$ $\downarrow$   $\text{CHAIR}_{i}$ $\downarrow$   Hall. $\downarrow$  Score $\uparrow$   Acc. $\uparrow$  F1 $\uparrow$
@@
DPO (RLAIF-V DataSet) LLaVA-1.5 7B 19.1 9.3 36.6 2.70 76.8 81.5
@@ 4.4 Analysis
Experiments on more advanced LVLMs.
Model Object-HalBench MMHal-Bench
LLaVA-Next(7B) 13.4 7.0 32.9 3.08
+DPO 11.9 6.8 33.6 2.99
+EMPO 5.3 3.7 29.2 3.18
Muffin (13B) 21.1 11.2 52.4 2.70
+DPO 12.7 6.3 43.9 2.82
+EMPO 7.0 4.7 40.8 2.90
Table 5: Experiments on more advanced LVLMs.
To validate the scalability and effectiveness of EMPO on more advanced LVLMs, we conduct experiments on two state-of-the-art models: LLaVA-Next 7B Liu et al. (2024b) and Muffin 13B Yu et al. (2023). As shown in Table 5, EMPO substantially reduces hallucination rates across both models, significantly outperforming their vanilla and DPO-trained counterparts. Specifically, for LLaVA-Next 7B, EMPO reduces 55.9% (11.2%) relative hallucination on Object-HalBench (MMHal-Bench). These results underscore EMPO’s consistent performance benefits across diverse LVLM architectures.
@@ 5 Conclusion
Acknowledgments
We thank all reviewers for their insightful comments and suggestions. This work is supported by the National Natural Science Foundation of China under Grants 62476188, the Natural Science Foundation of the Jiangsu Higher Education Institutions of China, Key Laboratory of New Generation Artificial Intelligence Technology & Its Interdisciplinary Applications (Southeast University), Ministry of Education, China.
@@ A.1 Quality Control
We analyze our entity-level editing strategy using CLIP-based semantic similarity, measuring the correspondence of an image region to both source and target entity labels before and after editing. This dual analysis objectively verifies whether the edit enhances the target entity’s presence while removing traces of the source. Using CLIP, we compute similarity scores. As shown in Figure 6(a), the results reveal a significant shift: post-edit, similarity to the target entity peaks near 0.8, while similarity to the source drops toward zero, reversing the pre-edit pattern. This demonstrates our method’s ability to both install target cues and erase source cues. This effect is robust across all five entity types (Figure 6(b)). In every category, the target’s median similarity rises post-edit as the source’s falls, confirming effective vision-language manipulation. "Person" and "car" categories benefit most, reflecting their strong visual signatures. Figure 6(c) quantifies this impact, showing an average similarity increase of 0.6 for the target entity and a corresponding decrease for the source. The minimal overlap between their score distributions indicates reliable substitution. Furthermore, Figure 6(d) shows that for similarity thresholds above 0.5, over 90% of edits succeed for the target entity, while the rate for the source approaches zero. This validates that our method produces semantically meaningful transformations. In summary, our CLIP-based analysis verifies that our pipeline enhances alignment with target entities while suppressing source entity evidence. This fine-grained control supports our claim that our dataset construction achieves clear, contrastive, and controllable entity-centric differences.
@@ A.2 Cost analysis
Having established robust quality safeguards, we now consider quantifying the computational requirements and monetary costs of each processing step. Extracting image entities costs about \$0.002 per entry; generating all 17,332 image entries on 8 A100 GPUs takes roughly 4–6 hours. Instruction-level entity extraction and rewriting add \$0.002 and \$0.003 per entry, respectively. In total, the labeling cost is \$0.007 per entry plus 8.3 seconds of compute time on a single A100. Constructing all 17,332 POVID entries cost approximately \$121 for GPT API fees and 40 hours of A100 computing resources (rented through Google Colab, about \$50). By contrast, manual annotation would cost approximately \$0.30 per entry, over 40× higher.
@@ Appendix B Hallucination Cause Analysis Experiment
Appendix C Extensive experiments on more comprehensive benchmarks.
To further evaluate EMPO’s performance, we conduct experiments on five additional benchmarks using the VLMEvalKit Duan et al. (2024) implementation. We use HallusionBench Guan et al. (2024) for hallucination evaluation and four others—MMMU Yue et al. (2024), MMBench Liu et al. (2024d), MMVet Yu et al. (2024d), and MIA-Bench Qian et al. (2024)—for general capabilities evaluation. As shown in Table 6, EMPO significantly mitigates hallucination, increasing the HallusionBench aA from 35.4% to 48.2%. It also consistently improves performance on MMBench, MMMU, MMVet, and MIA-Bench. These extensive results confirm that EMPO reduces hallucinations while maintaining or enhancing the model’s general proficiency.
Appendix D Prompt Appendix
The section describes the GPT4o-mini prompt for identifying entities, as well as the prompts for rewriting chosen and rejected instructions.
@@ Appendix D Prompt Appendix
Provide only the rewritten question without any additional explanation.
@@ 4. Output Format:
Appendix E Example Appendix
@@
Figure 7: Who in the picture is wearing trousers?